Lupo Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours.Multiple Choice О O $269.00 $189.00 О O $143.00 О O $67.25
Answer #1

Predetermined overhead rate = Variable overhead cost per machine hour + Fixed manufacturing overhead cost/Estimated machine hours

= 2 + 581,400/32,300

= 2 + 18

= $20 per machine hour

Overhead applied = Actual machine hours used x  Predetermined overhead rate

= 40 x 20

= $800

Job cost sheet

Direct material

630

Direct labor

1,260

Overhead applied

800

Total manufacturing cost

$2,690

Unit product cost = Total manufacturing cost/Number of units

= 2,690/10

= $269

First option is correct.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.